OpenStudy (anonymous):

y=2x^3+3x+4 find the inverse

OpenStudy (jdoe0001):

all you'd do is swap about the variables like so \(\bf y=2x^3+3x+4\qquad inverse\implies x=2y^3+3y+4\) and solve for "y"

OpenStudy (anonymous):

@majdishokri Are you still stuck?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Do you know how to complete the square?

OpenStudy (anonymous):

not good

OpenStudy (anonymous):

First divide by \(2\) because you want \(y^2\) on its own: \[ \frac x2=y^2+\frac 32y+2 \]

OpenStudy (anonymous):

but why we want to complete the square for this question

OpenStudy (anonymous):

Now you identify the middle coefficient is \(3/2\)

OpenStudy (anonymous):

So that we can solve for \(y\). You can also use the quadratic equation if you really want to.

hero (hero):

@wio, there's a \(y^3\) term

OpenStudy (anonymous):

y ^3 not y^2

OpenStudy (anonymous):

???????????????????????????????????????????????????

OpenStudy (anonymous):

Okay fine, first solve for \(y\) \[ y = \frac{x-4-2y^3}{3} \]And expand it: \[ y = \frac{x-4-2y^3}{3}=\frac{x-4-2\left(\frac{x-4-2y^3}{3}\right)^3}{3} \]

OpenStudy (anonymous):

What class is this, by the way?

OpenStudy (anonymous):

calculus 101

OpenStudy (anonymous):

Do you know how to differentiate?

OpenStudy (anonymous):

how u do that in the first step

OpenStudy (anonymous):

Maybe they want you to do implicit differentiation.

OpenStudy (anonymous):

this is a very complex question

OpenStudy (anonymous):

What have you learned recently?

OpenStudy (anonymous):

\[x=2y ^{3}+3y +4\] \[x-4=y(2y ^{2}+3)\] \[x-4\frac{ }{ ? }\]

OpenStudy (anonymous):

What concepts have you learned in your class recently?

OpenStudy (anonymous):

log

OpenStudy (anonymous):

????????

OpenStudy (anonymous):

log? Like logarithmic differentiation?

OpenStudy (anonymous):

yes